OpenStudy (anonymous):

A coyote had 10 pups in her third litter. It was twice as many as in her second litter, and her second litter had three more pups than her first. Find the number of pups in her first litter. How would you put that in a equation form? please i need help!

OpenStudy (anonymous):

A = first litter B = second litter C = third litter C = 10 C = 2B B=A+3 A=B-3=1/2C-3 now plug in numbers

OpenStudy (anonymous):

could you please write it as an equation, like a question, so then i could slove, please!

OpenStudy (anonymous):

A=1/2C-3 A=1/2(10)-3 A=5-3 A=2